### v1

# Ravenea: A Benchmark for Multimodal Retrieval-Augmented Visual Culture Understanding

Abstract
As vision-language models (VLMs) become increasingly integrated into daily life, the need for accurate visual culture understanding is becoming critical. Yet, these models frequently fall short in interpreting cultural nuances effectively. Prior work has demonstrated the effectiveness of retrieval-augmented generation (RAG) in enhancing cultural understanding in text-only settings, while its application in multimodal scenarios remains underexplored. To bridge this gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a new benchmark designed to advance visual culture understanding through retrieval, focusing on two tasks: culture-focused visual question answering (cVQA) and culture-informed image captioning (cIC). Ravenea extends existing datasets by integrating over 10,000 Wikipedia documents curated and ranked by human annotators. With Ravenea, we train and evaluate seven multimodal retrievers for each image query, and measure the downstream impact of retrieval-augmented inputs across fourteen state-of-the-art VLMs. Our results show that lightweight VLMs, when augmented with culture-aware retrieval, outperform their non-augmented counterparts (by at least $3.2\%$ absolute on cVQA and $6.2\%$ absolute on cIC). This highlights the value of retrieval-augmented methods and culturally inclusive benchmarks for multimodal understanding.
1 Introduction
Vision-language models (VLMs) are increasingly deployed in real-world applications, from education to assistive technologies di2024educational ; de2024vqask ; bates-etal-2025-template ; yao2025countllm , where understanding not only visual content but also the surrounding cultural context is crucial. Despite achieving impressive performance on general tasks yue2024mmmu ; lin2014microsoft ; an2025multimodality ; an2025generalized , VLMs often struggle to capture cultural nuances, such as traditions, symbols, and region-specific practices that require external, culturally grounded knowledge (nayak-etal-2024-benchmarking, ; liu2025culturevlm, ; khanuja-etal-2024-image, ; cao2024exploring, ). For example, as shown in Figure 1, a VLM may incorrectly identify the season of a festival scene as ‘Autumn’, overlooking that the image depicts Kyoto’s Gion Festival, which occurs in July and corresponds to ‘Summer’. A promising approach to address this limitation is the integration of external knowledge through retrieval-augmented generation (RAG) lewis2020retrieval , which has shown success in improving cultural awareness in language models (seo2025valuesrag, ; lertvittayakumjorn2025towards, ). However, prior work in the culture domain has predominantly been confined to text-only settings. Meanwhile, existing culture-related multimodal datasets primarily focus on evaluating VLM outputs on culturally oriented tasks, with limited emphasis on the integration of external cultural knowledge. As a result, the potential of RAG to improve multimodal cultural understanding remains underexplored.
Figure 1: Effectiveness of culture-aware RAG. Given a culturally grounded visual question, VLMs enhanced with culture-aware RAG—retrieving relevant Wikipedia documents—generate more accurate answers than their non-RAG counterparts. (see Section 5.2).
To bridge the gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a manually curated dataset designed to evaluate cultural understanding in VLMs with retrieval support. We construct Ravenea based on two existing datasets: CVQA (romero2025cvqa, ), which includes culturally relevant visual questions and corresponding answers, and CCUB (liu2023towards, ), offering culturally contextualized captions to foster inclusivity in text-to-image generation  [^1] We reuse the cultural captions from CCUB as ground-truth references for the inverse task, image-to-text generation, specifically for culture-aware image captioning.. For each instance drawn from the source datasets, we append a set of Wikipedia documents that have been human-ranked based on their cultural relevance to the associated image. This curation effort, designed to ensure broad cultural representation, contains data related to eight countries and spans eleven diverse categories, comprising more than 1,800 images and 10,000 human-ranked documents. Ravenea thus provides a retrieval-augmented benchmark for evaluating cultural sensitivity in multimodal retrievers, and further allows for assessing how well VLMs integrate and make use of retrieved cultural context. Specifically, we focus on two culturally grounded tasks: (i) culture-focused visual question answering (cVQA) and (ii) culture-informed image captioning (cIC). We select these two tasks for their centrality in assessing cultural understanding in VLMs—question answering tests context-aware reasoning, while captioning evaluates generation sensitivity to cultural nuances.
With Ravenea, we first train and evaluate seven multimodal retrievers that use both visual and textual inputs to retrieve Wikipedia documents for a given query image based on their cultural relevance. Then, we evaluate a diverse set of state-of-the-art (SOTA) VLMs, including GPT-4.1 achiam2023gpt , LLaVA-OneVision-7B li2025llavaonevision , Pixtral agrawal2024pixtral , Phi-4 Multimodal abouelenin2025phi , the Gemma3 family team2025gemma , the Qwen2.5-VL family bai2025qwen2 , the InternVL3 family zhu2025internvl3 , and the Deepseek-VL2 family wu2024deepseek , each with and without multimodal retrieval, to assess the impact of retrieval augmentation on cultural understanding. Our dataset provides a testbed for assessing the cultural relevance capabilities of multimodal retrievers and the effectiveness of VLMs in consuming and using such retrieved cultural context.
Our contributions and key findings include:
  • Ravenea benchmark: We introduce Ravenea, the first benchmark aimed at evaluating VLMs and multimodal retrieval in leveraging external knowledge for visual culture understanding. It comprises a diverse, large-scale collection of human-curated, culturally related documents linked to images from eight countries across eleven categories, enabling evaluation on two tasks: culture-focused visual question answering and culture-informed image captioning. (Section 3)
  • Cultural grounding annotations enhance multi-modal retrieval: We evaluate seven retrievers that integrate visual and textual cues to retrieve culturally relevant documents. We find fine-tuning retrievers on culture-targeted annotations leads to marked gains in retrieval accuracy, highlighting the value of explicit cultural supervision. (Section 6.3)
  • Benefits of culture-aware retrieval: Culture-aware retrieval boosts task performance across VLMs, with lightweight models showing the greatest improvement. This suggests that such retrieval can seamlessly integrate into downstream VLM tasks, enhancing their performance. (Section 5.2)
  • Cross-cultural variation: Evaluation across eight countries reveals that VLMs exhibit distinct cultural preferences, with each model favoring different regional contexts—suggesting model-specific cultural biases. (Section 6.2)
Figure 2: Ravenea: A Multimodal Retrieval-Augmented Visual culturE uNdErstAnding dataset. Left: Examples of cVQA and cIC tasks. Middle: Geographic and categorical distribution of cultural references. Right: Performance comparison of 14 VLMs, evaluated with and without integration of our culture-aware retriever. Here, CaCLIP="culture-aware CLIP-L/14@224px ".
2 Related Work
Retrieval augmentation for cultural understanding. Retrieval augmentation has demonstrated significant efficacy for culture-related NLP tasks (conia-etal-2024-towards, ; hu-etal-2024-bridging, ; fung-etal-2023-normsage, ). Prior work uses sources like the World Values Survey for cultural question answering seo2025valuesrag ; yadav-etal-2025-beyond , or retrieves web and knowledge base content to enhance cultural contextualization lertvittayakumjorn2025towards . In multimodal settings, however, cultural retrieval remains underexplored. While some approaches retrieve culture-relevant images to fine-tune VLMs liu2025culturevlm , they require additional training. In contrast, our method introduces a plug-and-play retrieval system that enriches cultural grounding at inference time without modifying the base model.
Vision-language culture datasets. Several recent studies examine cultural understanding in VLMs through multicultural VQA liu-etal-2021-visually ; nayak-etal-2024-benchmarking , cuisine recognition winata-etal-2025-worldcuisines ; li-etal-2024-foodieqa , and concept-based image retrieval bhatia-etal-2024-local . Others benchmark cultural entity recognition using Wikipedia-based prompts nikandrou-etal-2025-crope or curate culture-specific image sets for value-based tasks yadav-etal-2025-beyond . While these efforts highlight current VLM limitations, they often rely on static, manually curated datasets. A natural extension to address these deficiencies is through culture-relevant multimodal retrieval—an area lacking dedicated datasets. Our work specifically aims to fill this critical research gap.
3 Ravenea Dataset
We construct our dataset by building culturally relevant document lists for each image in two existing culture-grounded datasets: CVQA romero2025cvqa , a widely used dataset for culture-focused VQA, and CCUB liu2023towards , a dataset designed to mitigate cultural bias in text-to-image generation with culturally contextualized captions. To ensure broad geographic and cultural coverage, we curate a diverse subset comprising images from seven countries in CVQA: China, Nigeria, Russia, Spain, Mexico, India, and Indonesia, and all five countries in CCUB: China, Korea, India, Mexico, and Nigeria (see Figure 2). After that, we separate the dataset construction process into three critical stages shown in Figure 3.
3.1 Dataset Construction
Data collection. The data collection process consists of two main steps: (i) culture-related captioning and (ii) document retrieval. For culture-related captioning, we generate culturally grounded captions for each image to facilitate more effective attachment of relevant documents. Since the CVQA lacks captions and the CCUB provides only brief descriptions, we employ GPT-4o to generate richer, culturally informative captions (see the prompt example in Table 10). For document retrieval, we first conduct a coarse filtering using the generated cultural captions as queries for a BM25 Robertson2009ThePR retriever to extract the semantically relevant documents from a large-scale corpus comprising over six million English Wikipedia documents  [^2] https://huggingface.co/datasets/wikimedia/wikipedia. To mitigate the impact of inaccurate captions and ensure precise document relevance, we then perform human annotation on the retrieved documents.
Figure 3: Ravenea construction pipeline. Left: A two-stage retrieval process to match each image with relevant documents. Middle: Decomposition of cultural relevance into three interpretable dimensions to improve human annotation. Right: Postprocessing methods for quality control.
Relevance annotation. Based on the initial BM25 retrieval results, we refine the cultural relevance label of retrieved documents via human annotation. For each image-caption pair, annotators are presented with the top 10 Wikipedia documents retrieved by BM25. They are asked to assess whether each document provides meaningful background or contextual information that is relevant to the culture described in the caption or the image (see Appendix K). Specifically, we decomposed cultural relevance into three interpretable and independently verifiable dimensions: Country association: Is the topic of the Wikipedia article associated with the same country as the image and its caption? Topic alignment: Does the topic of the Wikipedia article align with the semantic category of the image and its caption? Explicit visual representation: Is the topic of the Wikipedia article explicitly mentioned or visually represented in the image and its caption? Each dimension is framed as a binary (True / False) question to reduce ambiguity and improve annotation consistency. However, for the country association dimension (the first listed), we introduce an additional label, "Cannot be determined", to handle cases where this association is unclear from the annotator’s perspective. Additionally, annotators are also instructed to include the title and URL of any relevant Wikipedia article they believe is missing from the top-10 retrieved results. These manually suggested articles are treated as the most cultural references closely related to the given image (see details in Appendix G).
Quality control. To ensure the quality and consistency of our annotations, we implement several quality control methods. First of all, prior to the annotation process, all annotators are required to carefully review a detailed instruction file outlining the relevance criteria and annotation guidelines.
Table 1: Statistics of the Ravenea dataset. The dataset is constructed by curating existing sources and augmenting them with over 10,000 wiki-derived documents to broaden cultural knowledge coverage and enhance content diversity.
Dataset Images Documents Pairs Questions Captions
CVQA 1,213 8,319 12,130 2,331 -
CCUB 655 4,441 6,550 - 655
Ravenea 1,868 11,580 18,680 2,331 655
To ensure proper understanding of the guidelines, annotators are required to complete a mock annotation test and correctly answer all questions before proceeding with the actual annotation tasks (see Figure 14). We also perform an additional quality check on a subset of the dataset. Specifically, for each selected countries, we employ an additional local quality checker who is tasked with manually reviewing the annotations to verify their accuracy and adherence to the guidelines. The quality checker reviews a random sample of annotated items, focusing on both the relevance labels and the justification behind any edge cases, such as borderline relevance or use of the “Cannot be determined” label. If inconsistencies or deviations from the annotation guidelines are identified, the affected samples are flagged for re-annotation. The overall acceptance rate from the meta quality checkers is 98.2%. The inter-annotator agreement (IAA) Cohen’s Kappa ($\kappa$artstein2017inter between the meta checker and annotator on the sampled annotations is 0.83.
Figure 4: Examples demonstrating the impact of CaCLIP Wikipedia retrieval integration on cVQA and cIC tasks using DeepseekVL2-Tiny. When augmented with culture-aware retrieval, the model exhibits enhanced sensitivity to cultural context.
3.2 Dataset Statistics
We present the statistics of our dataset in Table 1. The dataset comprises a total of 1,868 culturally diverse images, with approximately 65% originating from CVQA and the remaining 35% from CCUB. Each image is paired with a GPT-4o-generated cultural caption, as well as the top-10 ranked Wikipedia documents retrieved via a cultural relevance scoring pipeline (Figure 3), yielding a total of 18,680 image-document pairs (see illustrated examples in Figure 4). The collection spans eight countries and encompasses a broad spectrum of cultural domains, such as traditional attire, festivals, architecture, cuisine, and social practices (see more statistics in Appendix E).
4 Culture-aware Multimodal Retriever
Leveraging the Ravenea dataset, we train and evaluate seven multimodal retrievers to retrieve culturally relevant Wikipedia documents using both visual and textual inputs. We fine-tune five representative models—spanning both generative and discriminative paradigms—to optimize multimodal document retrieval. Performance is evaluated using standard retrieval metrics, including Mean Reciprocal Rank (MRR) shi2012climf , Precision@k (P@k) jarvelin2017ir , and Normalized Discounted Cumulative Gain (nDCG@k) wang2013theoretical , where $k\in\{1,3,5\}$. We integrate responses from three annotation questions per data point into a continuous scale ranging from $-3$ to $3$, where higher values indicate stronger cultural relevance. We fine-tune a VisualBERT-based li2019visualbert ; li-etal-2020-bert-vision reranker following standard BERT-style setups nogueira2019passage , and adapt two multimodal generators—VL-T5 cho2021unifying and LLaVA-OneVision-7B li2025llavaonevision —for end-to-end document retrieval chen2022corpusbrain ; yuan2024asking ; sun2023learning . To enhance cultural awareness in the contrastive retrieval, we introduce Culture-Aware Contrastive (CAC) learning, a supervised learning framework compatible with both CLIP and SigLIP architectures. We denote the culture-aware fine-tuned versions of CLIP-L/14@224px and SigLIP2-SO/14@384px using CAC as CaCLIP and CaSigLIP2, respectively.
4.1 Culture-aware Contrastive Learning
Given an image $I_{i}$ associated with $T$ textual descriptions $\{D_{i1},D_{i2},\dots,D_{iT}\}$, each document $D_{it}$ is annotated with a binary label $y_{it}\in\{0,1\}$, where $y_{it}=1$ indicates cultural relevance and $y_{it}=0$ indicates irrelevance. For each image–text pair $(I_{i},D_{it})$, we employ a shared vision-language encoder—such as CLIP—to obtain modality-specific representations: $\mathbf{E}_{I_{i}}=\mathcal{E}_{V}(I_{i})$ for the visual input and $\mathbf{E}_{D_{it}}=\mathcal{E}_{L}(D_{it})$ for the textual input. We then compute the cosine similarity score $s_{it}$ between $\mathbf{E}_{I_{i}}$ and each corresponding $\mathbf{E}_{D_{it}}$, resulting in a similarity vector $\mathbf{S}_{i}=[s_{i1},s_{i2},\dots,s_{iT}]$.
Culture-awareness classification now amounts to:
$\mathcal{L}_{\text{Culture Classify}}=-\frac{1}{B\cdot T}\sum_{i=1}^{B}\sum_{j=1}^{T}\left[y_{ij}\log\sigma(s_{ij})+(1-y_{ij})\log(1-\sigma(s_{ij}))\right],$ (1)
where B is the number of images; $\sigma(\cdot)$ denotes the sigmoid function.
To prioritize culturally relevant descriptions in the ranking, we apply a margin ranking loss between all pairs of descriptions with differing cultural relevance. For each image $I_{i}$, we compare all pairs $(D_{ij},D_{ik})$ such that $y_{ij}=1$ and $y_{ik}=0$, and encourage the model to assign a higher similarity score to the relevant description. The ranking loss is defined as:
$\mathcal{L}_{\text{Rank}}=\frac{1}{B}\sum_{i=1}^{B}\sum_{\begin{subarray}{c}j,k=1\\ y_{ij}=1,y_{ik}=0\end{subarray}}^{T}\max\left(0,\delta-(s_{ij}-s_{ik})\right),$ (2)
To mitigate the risk of overly similar positive text embeddings for the same image, we introduce a penalty that encourages intra-modal diversity among textual representations. We apply a diversity-promoting loss that forces the similarity between different text embeddings to be reduced while keeping each embedding highly similar to itself. Specifically, the penalty is formulated using an exponential function to emphasize the dissimilarity between embeddings:
$\mathcal{L}_{\text{Diversity}}\left(\mathbf{S}_{i}\right)=-\sum_{t=1}^{T}\log\left(\frac{\exp(s_{it})}{\sum_{j=1}^{T}\exp(s_{ij})}\right)$ (3)
Then we can get the culture-aware contrastive loss:
$\mathcal{L}_{\text{CAC}}=\frac{1}{3}\left(\mathcal{L}_{\text{Culture Classify}}+\mathcal{L}_{\text{Rank}}+\mathcal{L}_{\text{Diversity}}\right).$ (4)
Table 2: Performance with different retriever models. Fine-tuned contrastive models consistently outperform their frozen counterparts across tasks. Here, "CaSigLIP"=Culture-aware SigLIP2-SO/14@384px, and "CaCLIP"=Culture-aware CLIP-L/14@224px. Models in gray are frozen.
Method MRR $\uparrow$ P@1 $\uparrow$ P@3 $\uparrow$ P@5 $\uparrow$ nDCG@1 $\uparrow$ nDCG@3 $\uparrow$ nDCG@5 $\uparrow$ SigLIP2-SO/14@384px 66.71 50.42 40.14 34.25 58.85 62.43 67.50 CLIP-L/14@224px 70.76 54.58 43.47 36.58 62.54 67.25 72.43 VisualBERT 59.66 42.50 35.97 32.75 51.29 55.49 62.29 VL-T5 55.53 35.42 34.03 31.42 45.21 53.18 59.62 LLaVA-OneVision-7B 54.15 36.20 31.83 28.69 45.85 50.82 56.74 CaSigLIP2 (ours) 69.35 54.17 43.47 36.50 61.98 66.75 71.98 CaCLIP (ours) 78.34 65.42 49.44 39.50 72.25 75.22 79.32
4.2 Multimodal Retrieval Results
We perform a comprehensive evaluation of both frozen and fine-tuned retrievers, and present the results in Table 2. We find that fine-tuned models, particularly those based on contrastive learning, consistently outperform their frozen counterparts. For instance, CaCLIP achieves a substantial improvement in P@1, rising from 54.58% to 65.42%, and sets a new SOTA across all evaluation metrics. Although SigLIP2-SO/14@384px also benefits from fine-tuning, the performance gains are comparatively modest. In contrast, models such as LLaVA-OneVision-7B, VL-T5, and VisualBERT lag behind after fine-tuning, even underperforming relative to frozen baselines. This underperformance likely stems from the fact that models such as LLaVA-OneVision-7B and VisualBERT were originally pretrained for generative tasks with different objectives, whereas CLIP-L/14@224px and SigLIP2-SO/14@384px were explicitly trained for similarity-based alignment, providing them with a structural advantage in retrieval settings.
5 Multimodal Retrieval-augmented Visual Culture Understanding
We then evaluate the effectiveness of these retrievers with 14 SOTA VLMs, spanning a diverse set of architectures. We conduct experiments on two downstream tasks: cVQA and cIC, respectively.
5.1 Experimental Setup
Models.
We benchmark open and closed-weight SOTA VLMs on Ravenea, leveraging various retrievers against non-RAG baselines, assessing retrieval effectiveness across models of different sizes. The open-weight models include LLaVA-OneVision-7B li2025llavaonevision , Pixtral-12B agrawal2024pixtral , Phi-4 Multimodal-Instruct abouelenin2025phi , Gemma3-4B-Instruct and 27B-Instruct team2025gemma , Qwen2.5-VL-Instruct (3B, 7B, 72B  [^3] Due to computational constraints, we use the quantized version Qwen2.5-VL-Instruct-72B-AWQ.bai2025qwen2 , InternVL3 (2B, 8B, 38B) zhu2025internvl3 , and Deepseek-VL2 variants (Tiny and Base) wu2024deepseek . For the closed models, we adopt GPT-4.1 achiam2023gpt (accessed on 2025/04/14)  [^4] Knowledge cutoff: June 1, 2024; https://platform.openai.com/docs/models/gpt-4.1.
Evaluation metrics.
Table 3: Kendall’s $\tau$ rank correlation kendall1938new between automatic metrics and human judgments for the CCUB task. Statistically significant correlations ($p<0.05$) are marked with ✓. Our proposed metrics correlate stronger with human evaluation than the others.
Rouge-L lin-2004-rouge CIDER vedantam2015cider BERTScore Zhang*2020BERTScore: CLIPScore hessel-etal-2021-clipscore RegionScore (ours)
-0.172 ✗ -0.316 ✗ -0.011 ✗ 0.139 ✗ 0.442 ✓
For the cVQA task, we use accuracy as the primary evaluation metric, which measures the proportion of correctly predicted answers. For the cIC task, we employ several evaluation metrics including ROUGE-L lin-2004-rouge , CIDEr vedantam2015cider , BERTScore Zhang*2020BERTScore: , and CLIPScore hessel-etal-2021-clipscore , to assess the alignment between generated and reference captions across lexical, syntactic, and embedding-based levels. To further evaluate the cultural relevance and human-perceived quality, we conduct a human evaluation study. We employed four researchers to select the most accurate caption from 14 VLMs (see details in Appendix I and L), and find a significant mismatch between automatic metric scores and human judgments of cultural appropriateness (see Table 3). To bridge this gap, we further introduce RegionScore, a novel evaluation metric designed to quantify cultural grounding (see details in Appendix H and Table 9). It measures how well captions identify the correct country names tied to cultural elements, adding geographic and cultural specificity in image captioning.
5.2 Overall Performance
We present the main results in Table 4. The results demonstrate the efficacy of incorporating culture-aware retrieval augmentation. Employing fine-tuned retrievers yields substantial performance gains over both non-RAG and frozen retrievers baselines. Specifically, the CaCLIP achieves the highest average performance across both tasks, improves the accuracy for cVQA from 67.7% to 71.5%, and substantially improving the RegionScore for cIC from 40.2% to 58.1%. While CLIP-L/14@224px also offers improvements, fine-tuning consistently unlocks further potential. Furthermore, in the cVQA task, among all evaluated models, GPT-4.1 achieves the highest accuracy (86.8%) without RAG. Within the category of open-weight models, Qwen2.5-VL-72B leads with an accuracy of 81.0%. For lightweight models ($\leq$ 8B parameters), Qwen2.5-VL-7B achieves the best performance without RAG, reaching 67.7%. However, incorporating a CaCLIP significantly boosts performance, which enables InternVL3-8B to achieve 74.2% and outperform Qwen2.5-VL-7B by 0.6% with identical reranking. Notably, across multiple model families, augmenting the smallest variant with CaCLIP consistently elevates its performance to match or even exceed that of the next larger model tier. In the cIC task, with culture-aware contrastive learning, CaCLIP demonstrates substantial gains in identifying the culture in country-level of visual content, especially when built on top of VLMs with strong vision-language priors. It achieves the highest average RegionScore (58.1%) among the six reranking methods evaluated, with peak performance reaching 76.3% on the Gemma3-4B backbone. CaCLIP achieves leading the scores on 9 of 14 diverse suits of VLMs. This result underscores CaCLIP’s robustness and adaptability, particularly in culture-aware image captioning and retrieval tasks that demand fine-grained multimodal alignment.
Table 4: cVQA and cIC Performance w/ and w/o RAG. Models in gray are frozen retrievers. Results are colored as Best. VLMs augmented with finetuned retriever generally perform better.
Open Weights Closed Weights Retriever Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-38B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OneVision-7B GPT-4.1 $cVQA_{Accuracy\uparrow}$ W/O RAG 67.7 57.1 65.5 64.2 67.7 81.0 58.4 66.5 71.9 63.6 78.7 55.5 66.1 64.8 86.8 SigLIP2-SO/14@384px 67.2 58.7 55.8 63.9 67.7 77.1 62.3 68.4 71.9 60.7 71.9 62.6 69.4 67.1 82.6 CLIP-L/14@224px 70.6 64.2 56.1 67.7 73.2 79.0 68.1 71.0 76.1 66.1 74.5 69.4 71.0 69.7 81.9 Finetuned Models VisualBERT 67.1 58.7 59.4 66.5 68.4 77.1 61.3 66.1 71.9 60.7 73.6 63.2 66.5 63.9 82.3 VL-T5 65.8 56.5 61.3 62.9 69.4 77.1 60.3 66.8 70.3 55.8 74.2 61.0 61.0 61.9 82.6 LLaVA-OneVision-7B 67.8 60.0 61.3 64.8 71.3 80.7 61.3 64.8 73.9 62.3 73.6 62.6 67.4 63.9 81.3 CaSigLIP2 (ours) 69.8 62.6 58.7 67.1 72.3 77.7 65.8 72.6 75.8 64.2 75.5 64.2 70.7 69.4 82.6 CaCLIP (ours) 71.5 65.8 60.0 69.4 73.6 81.0 68.7 74.2 75.2 66.8 75.2 66.8 71.0 70.0 83.2 $cIC_{RegionScore\uparrow}$ W/O RAG 40.2 31.3 36.3 36.3 48.8 47.5 27.5 37.5 22.5 67.5 62.5 10.0 35.0 18.8 47.5 SigLIP2-SO/14@384px 52.5 50.0 56.3 36.3 57.5 53.8 58.8 56.3 61.3 70.0 66.3 30.0 53.8 41.3 53.8 CLIP-L/14@224px 55.4 56.3 63.8 41.3 66.3 60.0 61.3 62.5 67.5 70.0 70.0 31.3 55.0 47.5 58.8 Finetuned Models VisualBERT 56.3 58.8 63.8 41.3 58.8 62.5 61.3 61.3 66.3 71.3 70.0 30.0 58.8 45.0 62.5 VL-T5 55.4 60.0 65.0 41.3 62.5 66.3 62.5 58.8 67.5 71.3 68.8 30.0 61.3 40.0 61.3 LLaVA-OneVision-7B 56.0 60.0 65.0 40.0 60.0 62.5 61.3 62.5 66.3 71.3 71.3 28.8 57.5 47.5 60.0 CaSigLIP2 (ours) 56.3 57.5 65.0 36.3 61.3 63.8 58.8 61.3 67.5 71.3 70.0 27.5 57.5 46.3 65.0 CaCLIP (ours) 58.1 60.0 67.5 42.5 65.0 65.0 62.5 60.0 70.0 76.3 75.0 27.5 62.5 52.5 55.0
6 Analysis and Further Discussion
Culture-aware retrieval augmentation substantially benefits VLMs across both cVQA and cIC tasks, compared to their no-RAG counterparts. In this section, we explore the margin of the improvement, cultural preference and the effectiveness of cultural annotation.
6.1 Scaling Models Yields Diminishing or Negative Returns across Retrievers
Figure 5: Performance improvements for smallest and largest models per family with multimodal retrievers. Scaling models yields marginal gains with various retrievers, even negative effects in both cVQA and cIC tasks. "ACC." denotes accuracy; "R.S." refers to the RegionScore; "$\Delta$" represents the change incorporated with RAG compared to the non-RAG baseline.
In the cVQA task, within the same VLM family, performance differences between RAG and non-RAG approaches exhibit non-monotonic trends as model size scales. For all four model families, larger models show marginal or even negative returns from RAG integration. What’s more, sensitivity to RAG varies across model families. Notably, DeepSeek-VL2 demonstrates the most pronounced performance gap: the smallest model benefits from RAG with an average improvement of approximately +5%, whereas the largest model in the same family suffers a degradation of around -6% on average. In the cIC task, the effectiveness of RAG exhibits a consistent trend with respect to model scale within a given model family. Across all four model families evaluated, larger models tend to benefit less—or at most comparably—from the integration of RAG, suggesting diminishing returns at higher capacity. Among them, Gemma3 models show the smallest relative improvement, achieving approximately a +7% performance gain on average, whereas InternVL3 models yield the highest benefit, with performance gains reaching up to +30%.
6.2 Differences across Countries
We evaluate all models using CaCLIP across a range of countries for both tasks, as shown in Figure 6. In the cVQA setting, most VLMs exhibit substantially diminished performance on culture-specific questions regarding Nigeria and Indonesia, in contrast to their performance on questions under other national contexts. Interestingly, questions related to Spanish culture reveal high inter-model variance,
Figure 6: Performance of 14 VLMs with CaCLIP across different countries. Despite the integration of CaCLIP, disparities in model performance persist across countries.
with accuracy differentials reaching up to 35%, underscoring significant discrepancies in cultural representation across models. In the cIC task, VLMs consistently underperform on images and documents associated with Indian cultural contexts, while achieving the highest RegionScores on Korean culture-related inputs. Model performance on Indian culture is particularly volatile, indicating inconsistent cultural grounding across architectures. By comparison, Korean and Chinese cultural inputs yield more stable performance across models, suggesting entrenched model-specific preferences in cultural alignment (see more results in Appendix J).
6.3 Ablation Study on Annotation Questions
Figure 7: Ablation for different annotation questions. Combining all three culture-relevant annotation questions yields the best performance.
We further perform ablation studies across diverse combinations of annotation questions to assess their impact on downstream performance. Specifically, we evaluate 13 open-weight VLMs equipped with either CaSigLIP or CaCLIP, each trained on datasets constructed using varying subsets of culture-relevant annotations. From Figure 7, we can observe that leveraging all three questions (Q1 regarding country association; Q2 for topic alignment; Q3 for visual representation fidelity) yields the strongest performance on both cVQA and cIC tasks. For the cVQA task, we find Q1 provides the most significant benefit to CaSigLIP, whereas CaCLIP gains more from Q2. Among all pairwise combinations, the joint supervision from Q1 (country associations) and Q2 (topic alignment) proves slightly more effective than other pairs. In the cIC task, both CaSigLIP and CaCLIP achieve better performance improvements when trained with data derived from Q1, compared to other single-question sets. For pairwise combinations, CaCLIP benefits most from the Q1+Q3 combination, while CaSigLIP shows a clear preference for the Q2+Q3 setup.
7 Conclusion
We introduce Ravenea, a novel benchmark dataset designed to comprehensively evaluate the cultural sensitivity of diverse multimodal retrievers across 14 SOTA VLMs. The benchmark consists of culturally contextualized queries and image captions from eight countries, curated and paired with Wikipedia passages ranked by human annotators based on their cultural relevance. Our findings highlight the potential of RAG for visual cultural understanding, particularly when lightweight VLMs are enhanced with culturally-aware multimodal retrievers such as CaCLIP, which consistently outperform their non-augmented counterparts – with the caveat that the largest models in each VLM family often exhibit diminishing returns when integrated with RAG. Notably, the inclusion of culture-sensitive questions during data annotation significantly improves effectiveness of multimodal retrievers and enhances performance in downstream tasks.
8 Acknowledgment
We would like to thank all the annotators for their work. Special thanks to Nico Lang, Peter Ebert Christensen, Zhaochong An, Stella Frank, Srishti Yadav, for providing helpful research advice. Serge Belongie and Jiaang Li are supported by the Pioneer Centre for AI, DNRF grant number P1. Ivan Vulić is supported by a personal Royal Society University Research Fellowship ‘Inclusive and Sustainable Language Technology for a Truly Multilingual World’ (no 221137). Wenyan Li is supported by the Lundbeck Foundation (BrainDrugs grant: R279-2018-1145).
Appendix A Limitations
While Ravenea establishes a solid foundation for advancing the study of visual culture understanding with retrieval augmentation, it has three limitations that warrant future attention. First, due to budgetary constraints, the dataset’s cultural scope is currently limited to eight countries and eleven categories. Although this selection introduces meaningful diversity, it does not comprehensively represent the global spectrum of cultural perspectives, particularly those of underrepresented or marginalized communities. Second, our use of Wikipedia as the primary external knowledge source introduces inherent biases and may lack the depth, plurality, and contextual richness necessary for nuanced cultural interpretation. Finally, due to resource limitations, we were unable to include certain proprietary VLMs that require paid APIs, such as Gemini 2.5 Pro [52] and Claude Opus 3.7 [53]. We hypothesize that their performance would be comparable to GPT-4.1, which was included in our evaluation, but this remains an open empirical question.
Appendix B Future Directions
Our work opens several avenues for advancing visual culture understanding in multimodal models. First, expanding Ravenea to include more countries, cultural categories, and diverse knowledge sources, beyond Wikipedia, would improve coverage and reduce institutional bias. Second, future benchmarks could include richer tasks beyond cVQA and cIC, such as culture-grounded object recognition, historical retrieval, and symbolic interpretation, to better capture cultural semantics. Third, our results suggest a need for culturally-aware evaluation metrics, particularly for text generation. The limited effectiveness of retrieval augmentation in larger models also warrants further study, especially regarding how cultural knowledge is integrated and utilized. Together, these directions aim to support the development of more culturally-sensitive and globally robust vision-language models.
Appendix C Ethics Statement
This work focuses on improving the cultural awareness of VLMs through retrieval-augmented methods. All data used in the construction of the Ravenea benchmark were sourced from publicly available datasets and Wikipedia, a community-curated open-access knowledge base. To protect individual privacy, we apply automated face detection and blur all identifiable faces in images prior to release. To mitigate cultural bias and ensure broad representation, the benchmark includes images and documents spanning eight countries and eleven cultural domains, curated and annotated by a diverse group of annotators. We provide detailed documentation of the annotation process and guidelines to support transparency and reproducibility. While enhancing cultural understanding is a central goal of this work, we acknowledge that culture is inherently complex, dynamic, and context-dependent. Consequently, the benchmark cannot capture the full richness of any cultural context. Finally, this work does not involve any personally identifiable information (PII), biometric data, or sensitive attributes. All human annotators were compensated fairly, and data annotation adhered to ethical guidelines for responsible research.
Appendix D Experimental Details for Multi-modal Retrievers
Dataset.
We integrate responses from three annotation questions per data point into a continuous scale ranging from $-3$ to $3$, where higher values indicate stronger cultural relevance. To ensure fair evaluation across regions, we adopt the train-validation-test split strategy from class-imbalance loss [54], yielding an approximate 85-5-10 data split. This setup guarantees that validation and test sets contain a balanced set of images per country (Figure 8), thereby mitigating evaluation bias from skewed geographic distributions. During training, we employ random cropping for images as a sample-level augmentation to normalize the distribution of training instances across countries, further mitigating region-specific sampling biases.
Figure 8: Data distributions across eight countries. Restrict the dataset to the images paired with at least one document annotated by human raters as culturally relevant.
Table 5: Hyperparameters for finetuning on five models.
Hyperparameters VisualBERT VL-T5 LLaVA-OneVision-7B CLIP-L/14@224px SigLIP2-SO/14@384px
batch size 32 128 4 64 64
lr 2e-4 1e-4 1e-4 1e-5 1e-5
lr warmup ratio - 0.1 - - -
weight decay - - - - -
Max Epoch 100 20 10 50 50
Patience 10 10 3 5 5
early stopping Yes
optimizer AdamW
Using LORA - - Yes - -
Hyperparameters.
In this work, we adopt different sets of hyperparameters as VisualBERT, VL-T5, LLaVA-OneVision-7B, CLIP-L/14@224px, SigLIP2-SO/14@384px. For VisualBERT, VL-T5, and LLaVA-OneVision-7B we follow the setting in [44, 45, 46]. We show the training hyperparameters in rerankering experiments for all models in Table 5. All experiments are conducted using a maximum of 2 Nvidia H100 GPUs.
Appendix E Data Statistics
Table 6: Statistics of images in each country.
India Mexico Indonesia China Nigeria Korea Spain Russia
408 349 309 303 223 151 142 77
Table 7: Statistics of images in each category.
Architecture Cuisine History Art Daily Life Companies Sports & Recreation Transportation Religion Nature Tools
403 402 278 275 185 80 73 68 52 31 21
We then apply cryptographic hashing (SHA-256) to identify and remove duplicate images, resulting in a cleaner and more distinct cultural image set. Consequently, Ravenea comprises images collected from eight countries across four continents, spanning eleven distinct categories. The distribution of images by country and by category is detailed in Tables 6 and 7, respectively.
Appendix F Evaluation of VLMs
For closed-weight model, GPT-4.1, we directly call the corresponding API. For open-source models, we use vllm [55]. During the evaluation, to ensure the stability of the results, we set the temperature parameter to 0.0 and the maximum output length to 256. All open-weight models are listed in Table 8. To avoid the impact of the length the retrieved content, we use the first 256 words in the top-1 Wikipedia document.
Table 8: Model details: Hugging Face model names.
Model Hugging Face Model Name
LLaVA-OneVision-7B [19] llava-hf/llava-onevision-qwen2-7b-ov-hf
Phi-4-Multimodal [21] microsoft/Phi-4-multimodal-instruct
Pixtral [20] mistral-community/pixtral-12b
Qwen2.5VL family [23]
Qwen/Qwen2.5-VL-3B-Instruct
Qwen/Qwen2.5-VL-7B-Instruct
Qwen/Qwen2.5-VL-72B-Instruct-AWQ
DeepSeek-VL2 family [25]
deepseek-ai/deepseek-vl2
deepseek-ai/deepseek-vl2-tiny
InternVL3 family [24]
OpenGVLab/InternVL3-2B
OpenGVLab/InternVL3-8B
OpenGVLab/InternVL3-38B
Gemma3 family [22]
google/gemma-3-4b-it
google/gemma-3-27b-it
Appendix G Annotation Details
Based on the initial BM25 retrieval results, we refine the cultural relevance label of retrieved documents via human annotation. We found that directly asking annotators to rate overall cultural relevance on a continuous scale (e.g., 0–10) led to unreliable and inconsistent labels. This difficulty arises from several factors: (1) the semantic meaning of intermediate scores is ambiguous, (2) annotators tended to overemphasize a few salient visual elements [56], and (3) small numerical differences (e.g., between 5 and 6) often fail to reflect meaningful distinctions, especially given the cognitive load of processing lengthy Wikipedia documents, resulting in intra-annotator variance even on repeated examples. Instead, given an image–caption–document triplet, we decomposed cultural relevance into three interpretable and independently verifiable dimensions: country association, topic alignment, and explicit visual representation.
Prior to the annotation process, all annotators are required to carefully review a detailed instruction file outlining the relevance criteria and annotation guidelines. To ensure proper understanding of the guidelines, annotators are required to complete a mock annotation test and correctly answer all questions before proceeding with the actual annotation tasks. For each image-caption pair, annotators are presented with the top 10 Wikipedia documents retrieved by BM25. They are asked to assess whether each article provides meaningful background or contextual information that is directly relevant to the cultural elements described in the caption or depicted in the image.
Appendix H A New Metric: RegionScore
To evaluate the extent to which captions reference specific geopolitical regions, we introduce a RegionScore. This metric measures whether a caption contains explicit references to a country or its common demonyms. $N$ be the total number of samples. $\text{PRED}_{i}$ denote predicted captions for the $i$-th sample. $C_{i}$ be the country name associated with the $i$-th sample. $\text{Adj}(C_{i})$ denote the set of adjectives or demonyms associated with $C_{i}$. $\mathcal{T}_{i}=\{C_{i}\}\cup\text{Adj}(C_{i})$ denote the set of region-related terms for the $i$-th sample. We define binary indicators for each sample:
$\delta^{\text{pred}}_{i}=\begin{cases}1&\text{if any }t\in\mathcal{T}_{i}\text{ appears in }\text{PRED}_{i}\\ 0&\text{otherwise}\end{cases}$ (5)
The RegionScore for predicted captions are then computed as:
$\text{RegionScore}_{\text{PRED}}=\frac{1}{N}\sum_{i=1}^{N}\delta^{\text{pred}}_{i}$ (6)
These scores reflect the proportion of captions that include explicit regional identifiers. A higher score indicates stronger region-awareness in the captioning. For the ground truth in CCUB dataset, the $\text{RegionScore}_{\text{GT}}=99\%$.
Appendix I Details of Human Evaluation in the cIC Task
We randomly sample 10 images and generate captions using 14 vision-language models (VLMs) under three configurations: CaCLIP, CLIP, and a no-retrieval baseline, yielding 420 captions in total. Four expert annotators participated in the evaluation. For each image, they were presented with caption triplets—one per retrieval configuration—and asked to assess approximately 35 such triplets each. Annotators are instructed to select the caption that most accurately and appropriately reflects the cultural context depicted. To evaluate annotation consistency, we randomly sample 30 triplets and assign them to a fifth expert annotator. Inter-annotator agreement (IAA) between this annotator and the original four, measured using Cohen’s $\kappa$, is 0.595.
Appendix J Additional Results
Figure 9: Performance of the 14 VLMs equipped with CaCLIP in cVQA task.
As illustrated in Figures 9 and 10, the models exhibit diverse cultural preferences. Notably, most models achieve relatively stronger performance on Chinese and Indian cultural contexts in the cVQA task, and on Chinese and Korean contexts in the cIC task.
Figure 10: Performance of the 14 VLMs equipped with CaCLIP in cIC task.
We report the CIDEr, ROUGE-L, BERTScore, and CLIPScore metrics for 14 VLMs on the cIC task in Table 9.
Table 9: Four metrics comparison of RAG and Non-RAG methods for CCUB task.
Open Weights Closed Weights Method Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-38B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OneVision-7B GPT-4.1 $CCUB_{Rouge-L}$ W/O RAG 18.3 18.7 18.2 21.8 18.2 16.5 17.8 19.6 19.8 16.7 15.9 20.6 15.4 21.4 19.6 Frozen Models SigLIP2-SO/14@384px 18.0 20.0 19.1 23.8 17.2 14.9 19.2 17.4 17.5 15.1 14.9 23.1 14.0 23.6 17.0 CLIP-L/14@224px 18.1 20.9 18.5 23.0 16.8 14.5 18.3 17.2 17.5 15.1 15.8 22.4 14.4 24.7 16.1 Finetuned Models VisualBERT 18.6 19.5 18.3 22.5 16.8 14.8 17.4 17.4 16.9 15.0 16.0 22.5 14.8 24.3 19.0 VL-T5 17.6 19.0 18.1 23.1 17.9 14.4 18.2 17.2 16.9 14.7 15.4 20.8 13.8 23.5 17.4 LLaVA-OneVision-7B 17.3 18.4 17.7 23.0 17.2 14.5 17.6 16.4 17.1 14.6 15.0 20.9 13.1 22.8 17.3 SigLIP2-SO/14@384px 18.4 19.3 19.2 24.5 18.5 14.8 18.4 18.2 17.2 15.2 16.4 22.2 14.3 23.5 18.9 CLIP-L/14@224px 18.1 19.6 18.6 23.9 17.1 14.7 18.0 17.3 17.7 16.0 16.3 21.5 13.7 24.2 17.0 $CCUB_{CIDER}$ W/O RAG 28.1 30.0 27.2 49.2 22.9 16.9 10.5 18.6 35.8 11.2 10.1 49.1 16.7 47.4 34.4 Frozen Models SigLIP2-SO/14@384px 25.2 39.2 20.5 58.5 16.9 7.6 12.4 7.1 6.3 4.7 6.1 62.0 10.1 51.4 17.2 CLIP-L/14@224px 23.8 40.9 18.8 59.6 13.0 6.2 9.6 7.8 4.8 4.7 6.6 56.3 11.1 49.7 14.6 Finetuned Models VisualBERT 26.1 40.2 24.0 60.9 16.3 8.2 10.0 7.5 5.1 6.4 10.2 51.1 15.1 49.5 24.4 VL-T5 24.5 37.0 20.6 51.1 16.0 7.3 10.2 5.0 4.0 6.8 7.5 52.5 10.3 48.4 21.6 LLaVA-OneVision-7B 21.7 38.8 18.5 62.6 13.2 4.8 8.4 5.7 3.4 5.7 4.5 48.9 11.4 42.2 17.6 SigLIP2-SO/14@384px 28.2 40.2 22.8 64.7 17.8 8.4 9.7 5.8 4.4 4.7 7.7 62.3 13.2 54.8 26.5 CLIP-L/14@224px 24.6 35.4 22.1 61.5 15.7 8.9 9.9 6.5 5.8 5.8 7.8 53.1 11.2 51.7 18.3 $CCUB_{BERTScore}$ W/O RAG 54.7 55.4 55.6 55.6 55.7 54.2 52.5 54.5 55.4 54.1 54.8 53.5 52.7 56.6 56.8 Frozen Models SigLIP2-SO/14@384px 54.9 57.2 56.4 57.9 55.2 53.0 55.1 54.8 55.0 52.4 54.0 56.9 52.6 58.7 54.8 CLIP-L/14@224px 54.8 57.6 56.2 57.7 54.8 52.4 54.8 55.1 55.3 52.6 53.7 55.9 52.5 59.4 54.6 Finetuned Models VisualBERT 54.8 57.1 56.5 57.6 54.9 52.8 54.6 54.9 54.4 52.1 53.8 55.7 52.8 58.5 56.2 VL-T5 54.2 56.7 56.0 58.1 54.9 52.6 53.9 54.0 54.4 52.0 53.4 54.4 52.1 57.8 55.7 LLaVA-OneVision-7B 54.2 56.6 55.5 58.2 54.5 52.1 53.9 54.2 54.6 51.6 53.3 55.0 51.8 58.3 54.9 SigLIP2-SO/14@384px 55.2 57.4 57.1 57.9 55.7 53.2 54.7 55.2 54.9 52.5 54.2 56.8 52.6 58.7 56.7 CLIP-L/14@224px 55.0 57.0 56.5 58.2 54.8 53.2 54.6 55.3 55.3 52.9 54.2 56.0 52.3 59.3 55.4 $CCUB_{CLIPScore}$ W/O RAG 19.1 19.1 18.8 19.6 18.3 18.2 18.4 18.7 19.2 18.4 18.4 20.2 18.5 20.2 18.7 Frozen Models SigLIP2-SO/14@384px 19.1 19.3 19.0 19.9 18.9 18.8 19.1 18.8 18.7 18.6 18.5 20.5 18.8 20.0 18.6 CLIP-L/14@224px 19.0 19.2 19.0 20.0 19.0 18.5 19.2 18.8 18.6 18.8 18.4 20.4 18.5 19.6 18.3 Finetuned Models VisualBERT 19.1 19.1 18.9 20.2 18.9 18.9 19.1 18.9 18.8 18.5 18.6 20.3 18.7 19.9 18.5 VL-T5 19.1 19.4 19.1 19.9 19.0 18.8 19.0 19.1 18.7 18.5 18.7 20.4 18.7 19.9 18.4 LLaVA-OneVision-7B 19.1 19.3 18.9 20.1 18.9 18.7 19.3 18.9 18.6 18.5 18.6 20.4 18.8 19.7 18.6 SigLIP2-SO/14@384px 19.1 19.2 18.9 20.1 18.8 18.6 19.1 19.0 18.5 18.6 18.5 20.4 18.8 19.8 18.5 CLIP-L/14@224px 19.0 19.2 18.9 20.0 19.1 18.6 19.2 18.9 18.7 18.6 18.4 20.1 18.8 19.7 18.6
Figure 11: Average improvement across 14 VLMs in different countries with 4 retrievers for cVQA task.
As shown in Figures 12 and 11, VLMs demonstrate varying degrees of performance shifts across cultural contexts in both the cIC and cVQA tasks. In the cIC setting, most fine-tuned retrievers yield noticeable improvements over their original counterparts, indicating the effectiveness of retrieval adaptation. Performance on the cVQA task reveals more nuanced outcomes. While fine-tuned retrievers generally exhibit large performance improvement compared to non-RAG baselines, certain countries, such as Spain and Indonesia, experience exhibit diminishing returns. These discrepancies may stem from the limited presence of culturally representative visual content in the training data. Although targeted image augmentation strategies were employed to alleviate this imbalance, the results suggest that data distribution remains a significant bottleneck. Understanding and addressing such cross-cultural performance disparities in multimodal tasks like cVQA will be an important direction for future work.
Figure 12: Average improvement across 14 VLMs in different countries with 4 retrievers in cIC task.
Appendix K Human Annotation & Evaluation Interface
The annotation interface are shown in Figure 14 and 15. The interface for human evaluation in cIC task is shown in Figure 13.
Figure 13: Human evaluation interface for cIC task.
Figure 14: Human annotation instructions.
(a) Annotation interface (1).
(b) Annotation interface (2).
Figure 15: Human annotation interfaces.
Appendix L Correlation between Automatic Metrics and Human Judgments
To assess the correlation between human preferences and automatic evaluation metrics, we compute Kendall’s Tau rank correlation. Human annotations are segmented according to the output chunks produced by corresponding 14 VLMs with each retriever variant (CaCLIP-based, CLIP-based, and non-RAG). Within each segment, we calculate the selection winning ratio for each retrieval method, yielding a human preference vector formed by concatenating these ratios across all evaluation instances. For the automatic evaluation, we extract BERTScore, CIDEr, ROUGE-L, and CLIPScore for each corresponding retrieval variant and VLM. These scores are similarly concatenated into a metric-based vector. Finally, we compute Kendall’s Tau between the human and metric vectors to quantify the consistency between automatic rankings and human judgments.
Appendix M Prompts
The prompts that we used to collect the documents and two downstream tasks were designed to ensure consistency across different models. One prompt example for cultural caption generation is shown in Table 10. The prompts for both with and without retrieval augmentation for VLMs are shown in Table 12 and 11.
Table 10: Prompt example for generating captions used to retrieve the Wiki documents.
Prompt for data collection GPT-4o SYSTEM: Default
USER:
Generate a culture related caption given the image with around 2 sentences. Please include the name of the thing shown in the picture within the caption if it’s strongly related to the culture. Name the three most culturally relevant entities in the image and attach the name at the end. Follow the format of: caption. entity name1; entity name2; entity name3.
Table 11: Prompt examples without RAG. Multimodal prompt samples with interleaved image are shown for both CVQA and CCUB tasks.
CVQA task CCUB task
SYSTEM:
You are a helpful assistant.
USER:
Answer the following multiple choice question. The last line of your response must be of the following format: Ánswer: \$LETTER(́without quotes) where LETTER is one of ABCD.
Question: Which city is famous for such artwork?
A) Zhejiang Yiwu
B) Jingdezhen, Jiangxi
C) Datong, Shanxi
D) Zibo, Shandong
SYSTEM:
You are a helpful assistant.
USER:
Write a concise, one-sentence caption for the given image. The generated caption must contain the visual content and culturally relevant elements of the image. Avoid explicit references to the image itself (e.g., "This image shows...", "Pictured here is...", "In this photograph..."). Do not generate multiple options.
Table 12: Prompt examples with RAG. Multimodal prompt samples with interleaved image are shown for both CVQA and CCUB tasks.
CVQA task CCUB task
SYSTEM:
You are a helpful assistant.
USER:
Answer the following multiple choice question. The last line of your response must be of the following format: Ánswer: \$LETTER(́without quotes) where LETTER is one of ABCD.
The scope of the question is strictly limited to the given image. However, please analyze and incorporate information from both the image and the following document to answer the question.
Document:
Buildings and structures Buildings about 800 – Borobudur temple in Java completed. 802 Haeinsa of Korea, is constructed. Palace of Charlemagne in Aachen, Carolingian Empire completed (begun about 790). The Palatine Chapel still stands. At Oviedo in the Kingdom of Asturias Cámara Santa constructed. First reconstruction of Oviedo Cathedral begun by Tioda. 815 – Second Temple of Somnath built in the Pratihara Empire, India. 816 – Reims Cathedral begun. 810s – Chapel of San Zeno in Santa Prassede, Rome decorated. 818 – Old Cologne Cathedral built....
Question: What is the name of the square where the cathedral and the statue of the image are located?
A) Riego Square
B) Trascorrales Square
C) The square of Alfonso II the Chaste
D) The Fontan square
SYSTEM:
You are a helpful assistant.
USER:
Write a concise, one-sentence caption for the given image. The generated caption must contain the visual content and culturally relevant elements of the image. Avoid explicit references to the image itself (e.g., "This image shows...", "Pictured here is...", "In this photograph..."). Do not generate multiple options. Please consider the following context:
Architecture of Nigeria was historically influenced by environmental conditions as well as social and cultural factors. The coming of missionaries and political changes brought about by colonialism precipitated a change in architectural style and utility of buildings. A Gothic revival style was adopted for early churches built in the colony of Lagos. A one or two storey timber house building made with pre-fabricated material components and designed with the influence of classic antiquity styles served as mission house for the missionaries...